# A Hierarchical Approach to Quantum Many-Body Systems in Structured Environments

### v1

@@ Appendix B Derivation of the BBGKY-HEOM equations
Appendix C Multipolar gauge
@@
which corresponds straight line from the origin to the point $x^{\prime}$. This particula choice is motivated by the construction of the multipolar gauge [74]. We define a gauge function $\chi_{\mu}$ as
@@
On the other hand, in the multipolar gauge ($\mu=1$) the transition dipole of the system $d_{i}$ couples to the field $\Pi_{1}$. Physically this means that the system couples to the field which is the differnce of the transverse electric field generated bu the transverse vector potential and the polarization field
@@
Appendix D Derivation of the HEOM
We sketch here a derivation of the HEOM equations. We assume that the system is coupled to the environment via operator $q$. The response of the system to the environment is given by the bath correlation function $\alpha(t)$. For HEOM we need that the bath correlation function is of the form
$\alpha(t)=Ge^{-Wt},$ (42)
where $G$ and $W$ may be complex numbers. The reduced dynamics of the system can be written in the path integral form
We define auxiliary states
$\rho_{t}^{n,m}(q,q^{\prime})=$ $=\int\mathrm{d}q_{0}\int\mathrm{d}q_{0}^{\prime}\int\mathcal{D}q\int\mathcal{D}q^{\prime}\,$
$\dot{\rho}_{t}^{n,m}=$ $-i[H,\rho_{t}^{n,m}]-(nW+mW^{*})\rho_{t}^{n,m}$
$+[\rho_{t}^{n+1,m},q]+[q,\rho_{t}^{n,m+1}]$
$+nGq\rho_{t}^{n+1,m}+mG^{*}\rho_{t}^{n,m+1}q.$ (46)
The equations presented in the main text are a generalization of the above equations in two respects. First, we allow for multiple exponential terms in the expansion for the bath correlation function, and second we allow for the case $q\neq q^{\dagger}$.
Appendix E Numerical details
@@
Numerical simulations were performed with the python packages numpy[76], scipy[77] and qutip[78]. The BBGKY-HEOM equation (11) with the approximation for the three-body reduced density matrix given in Eq. (12) were solved with the scipy implementation of the explicit Runge-Kutta method of order 5(4), with an absolute tolerance of $101e-$ 10 and dynamical timestepping. We checked for the convergence with respect to the hierarchy depth, which justifies a hierarchy depth of 5 for the Tavis-Cummings model and 3 for the Cavity-Fermi-Hubbard model. The benchmark simulations for the quantum master equation were performed using qutips mesolve function with default parameters.

### v2

@@ Appendix C Derivation of the BBGKY-HEOM equations
Appendix D Local baths
In this section we show how BBGKY-HEOM can be used to treat situations, where each particle couples to its own local bath. For the sake of readability we consider the case where the bath correlation function of each local bath is a single exponential, where, again, the generalization to multiple exponential terms is straight forward. We start out with the HEOM equation for the full N-particle state, where the coupling operators $L_{k}=S_{k}$ only act on particle $k$ and $H_{sys}=\sum_{i=1}^{N}H_{sys}^{i}$
$\begin{split}\dot{\rho}^{(\bm{n},\bm{m})}&=-i[H_{sys},\rho^{(\bm{n},\bm{m})}]-\left(\bm{w}\cdot\bm{n}+\bm{w}^{*}\cdot\bm{m}\right)\rho^{(\bm{n},\bm{m})}\\ &+\sum_{k=1}^{N}\bigg{(}G_{k}\left(n_{k}S_{k}\rho^{(\bm{n}-\bm{e}_{k},\bm{m})}+m_{k}\rho^{(\bm{n},\bm{m}-\bm{e}_{k})}S_{k}^{\dagger}\right)\\ &+\left[\rho^{(\bm{n}+\bm{e}_{k},\bm{m})},S_{k}^{\dagger}\right]+\left[S_{k},\rho^{(\bm{n},\bm{m}+\bm{e}_{k})}\right]\bigg{)},\end{split}$ (40)
Now after tracing out particle $N$, with $\rho_{1..N-1}^{(\bm{n},\bm{m})}=\operatorname{Tr}_{N}(\rho^{(\bm{n},\bm{m})})$ we are left with
$\begin{split}\dot{\rho}_{1..N-1}^{(\bm{n},\bm{m})})&=-i[\sum_{i=1}^{N-1}H_{sys}^{i},\rho_{1..N-1}^{(\bm{n},\bm{m})}]-\left(\bm{w}\cdot\bm{n}+\bm{w}^{*}\cdot\bm{m}\right)\rho_{1..N-1}^{(\bm{n},\bm{m})}\\ &+\sum_{k=1}^{N-1}\bigg{(}G_{k}n_{k}S_{k}\rho^{(\bm{n}-\bm{e}_{k},\bm{m})}_{1..N-1}+G_{k}^{*}m_{k}\rho_{1..N-1}^{(\bm{n},\bm{m}-\bm{e}_{k})}S_{k}^{\dagger}\\ &\quad\quad\quad+\left[\rho_{1..N-1}^{(\bm{n}+\bm{e}_{k},\bm{m})},S_{k}^{\dagger}\right]+\left[S_{k},\rho_{1..N-1}^{(\bm{n},\bm{m}+\bm{e}_{k})}\right]\bigg{)}\\ &+G_{N}n_{N}\operatorname{Tr}_{N}(S_{N}\rho^{(\bm{n}-\bm{e}_{N},\bm{m})})\\ &+G_{N}^{*}m_{N}\operatorname{Tr}_{N}(\rho^{(\bm{n},\bm{m}-\bm{e}_{N})}S_{N}^{\dagger})\Big{)}.\end{split}$ (41)
We see that the last two terms actually still depends on hierarchy states with $n_{N},m_{N}>0$. However, ultimately we are only interested in the physical density matrix $\operatorname{Tr}_{N}(\dot{\rho}^{(\bm{0},\bm{0})})$. We can see that the evolution of this physical density matrix only depends on auxiliary states with $m_{N}=n_{N}=0$, because the sum in line two and three only extends to $N-1$. Furthermore, the evolution of all auxiliary states with $m_{N}=n_{N}=0$ do in turn also not depend on auxiliary states with $m_{N},n_{N}>0$. Thus the physical density matrix is completely independent of this part of the hierarchy and it may as well be omitted. After omitting the last line in Eq. (41) we have thus eliminated the particle $N$ together with its local bath. Continuing to trace out all particles except 1 and 2 we are then left with
$\begin{split}\dot{F}_{12}^{(\bm{n},\bm{m})}&=-i[H_{sys}^{1}+H_{sys}^{2},F_{12}^{(\bm{n},\bm{m})}]-\left(\bm{w}\cdot\bm{n}+\bm{w}^{*}\cdot\bm{m}\right)F_{12}^{(\bm{n},\bm{m})}\\ &+\sum_{k=1}^{2}\bigg{(}G_{k}\left(n_{k}L_{k}F_{12}^{(\bm{n}-\bm{e}_{k},\bm{m})}+m_{k}F_{12}^{(\bm{n},\bm{m}-\bm{e}_{k})}L_{k}^{\dagger}\right)\\ &+\left[F_{12}^{(\bm{n}+\bm{e}_{k},\bm{m})},L_{k}^{\dagger}\right]+\left[L_{k},F_{12}^{(\bm{n},\bm{m}+\bm{e}_{k})}\right]\bigg{)}.\end{split}$ (42)
Appendix E Multipolar gauge
@@
which corresponds to a straight line from the origin to the point $x^{\prime}$. This particular choice is motivated by the construction of the multipolar gauge [99]. We define a gauge function $\chi_{\mu}$ as
@@
On the other hand, in the multipolar gauge ($\mu=1$) the transition dipole of the system $d_{i}$ couples to the field $\Pi_{1}$. Physically this means that the system couples to the field which is the difference of the transverse electric field generated bu the transverse vector potential and the polarization field
@@
Appendix F Derivation of the few emitter lasing Hamiltonian
F.1 Huang-Rhys factors from TDDFT
All (TD)DFT calculations have been performed with the ORCA code [101] using the PBE exchange-correlation potential [102] with Grimme D3 dispersion correction [103] and def2-TZVPD basis. We expect the (electronic) structure of methylene-blue (MB) molecules to depend heavily on the embedding in cucurbit[7]uril and in the presence of the nanoparticle-on-mirror set-up. For simplicity, we have assumed here the oxidized form (positively charged MB+) of a single MB molecules, which is typically associated with an intense blue color.
First, the ground-state geometry was converged, and electronic absorption spectra were calculated using the Casida-TDDFT approach (employing the Tamm-Dancoff approximation). Second, the geometry of the S1 and S2 states have been relaxed. We found the S1 state to be virtually dark for absorption and emission (with a Stokes shift of 0.269 eV) and have therefore decided to focus on the S2 transition (Stokes shift 0.119 eV) for subsequent discussion. We note that the selected level of theory is insufficient for direct comparison with experimental spectra as the encapsulation can be expected to play a major role. Using higher-level exchange-correlation potentials (B3LYP+D4) and a continuous polarization model for water flips the excitation order (S1 is then bright) but leads otherwise to minor changes in the excitation frequency (486.8 nm) and oscillator strength at a considerably increased cost. The Tamm-Dancoff approximation has, on the other hand, a notable impact. Full Casida-TDDFT shifts the bright absorption (S1 now) to 550.5 nm and future studies that aim at full theoretical consistency are therefore advised to avoid the Tamm-Dancoff approximation.
In the following we give a brief derivation of Eq. (III.2) and show how the coupling strengths and frequencies of the vibrational modes are obtained from the DFT calculations. In the harmonic approximation the (classical) Hamiltonian governing the dynamics of the nuclear coordinates $\bm{\eta}$, expressed as deviations from their respective equilibrium positions can be given in terms of the mass matrix $\mathbf{M}$ and the Hessian Matrix of the potential energy $\mathbf{K}$ (obtained from the DFT calculations) as
$\begin{split}H=&\frac{1}{2}\dot{\bm{\eta}}^{T}\textbf{M}\dot{\bm{\eta}}+\frac{1}{2}\bm{\eta}^{T}\textbf{K}\bm{\eta},\\ =&\frac{1}{2}\dot{\bm{\mu}}^{2}+\frac{1}{2}\bm{\mu}^{T}\textbf{M}^{-1/2}\textbf{K}\textbf{M}^{-1/2}\bm{\mu},\\ \bm{\mu}=&\textbf{M}^{1/2}\bm{\eta}.\end{split}$ (58)
The mass re-normalized vibrational eigenmodes $u_{k}$ and the corresponding frequencies $\omega_{l}$ are now found by means of diagonalization
$\left[\textbf{M}^{-1/2}\textbf{K}\textbf{M}^{-1/2}\right]\mathbf{u}_{k}=\omega_{k}^{2}\mathbf{u}_{k}.$ (59)
In the electronic groundstate the vibrations are now decoupled in the eigenbasis
$\begin{split}x_{k}=&\bm{\mu}\cdot\mathbf{u}_{k},\\ H=&\frac{1}{2}\sum_{k}\dot{x}_{k}^{2}+\omega_{k}^{2}x_{k}^{2}\end{split}$ (60)
and can be canonically quantized
$\hat{H}=\sum_{k}\frac{\hat{p}_{k}}{2}+\frac{\omega_{k}^{2}}{2}\hat{q}_{k}^{2}.$ (61)
In the double harmonic approximation the dynamics in electronically excited state are governed by the same frequencies but shifted equilibrium positions $\bm{\mu}\to\bm{\mu}-\bm{\Delta}$, where $\bm{\Delta}$ is the mass weighted nuclear displacement between absorption (at 479.5 nm) and emission (at 502.5 nm) $\bm{\Delta}=\textbf{M}^{1/2}(\textbf{R}_{S0}-\textbf{R}_{S2})$. Along the $k$-th eigenmode we then obtain a displacement of $\Delta_{k}=\bm{\Delta}\cdot\textbf{u}_{k}$ in the excited state, leading the the Hamiltonian $H_{mol}$ in Eq. (III.2). The Huang-Rhys factors defined as $S_{k}=\omega_{k}\Delta_{k}^{2}/2$ are thus obtained as
$S_{k}=$ $\frac{\omega_{k}}{2}((\textbf{R}_{S0}-\textbf{R}_{S2})\cdot\sqrt{\textbf{M}}\cdot\textbf{u}_{k})^{2}.$ (62)
F.2 Effective spectral density of intra-molecular vibrations
Eq. (III.2) describes an isolated molecule, where the vibrations form the stick spectrum shown in Fig. 7.
Figure 7: Squared coupling strengths of the vibrational modes (blue) as obtained from DFT and ohmic spectral density of the phonon environment (red).
To account for the environment in which the molecules are embedded we couple each molecular vibration to an (accoustic) phonon bath with an ohmic spectral density. It is important to note that the phonon bath will always relax the vibrational modes to their respective groundstate, which is shifted if the molecule is electronically excited. To ensure this, we model the coupling between the molecular vibrations and the environmental phonon modes with the coupling operator $b_{\lambda}-\Delta_{\lambda}\sigma_{+}\sigma_{-}$ [104]. The total Hamiltonian for a single molecule and the phonon bath with annihilation operators $d_{\xi}$ now reads
$\begin{split}H_{mol,env}=&\frac{\omega_{0}}{2}\sigma^{z}-\frac{1}{\sqrt{2}}\sum_{\lambda=1}^{107}\omega_{\lambda}^{vib}\sqrt{S_{\lambda}}\sigma_{+}\sigma_{-}(b_{\lambda}+b_{\lambda}^{\dagger})\\ &+\sum_{\lambda=1}^{107}\omega_{\lambda}^{vib}b_{\lambda}^{\dagger}b_{\lambda}+\sum_{\lambda,\xi}\omega_{\xi}d_{\lambda,\xi}^{\dagger}d_{\lambda,\xi}\\ &+\sum_{\lambda,\xi}g_{\xi}(b_{\lambda}-\Delta_{\lambda}\sigma_{+}\sigma_{-})d_{\lambda,\xi}^{\dagger}+\mathrm{h.c.},\\ J_{ohm}(\omega)=&\sum_{\xi}\lvert g_{\xi}\rvert^{2}\delta(\omega-\omega_{\xi}),\\ =&A\omega\exp(-\omega/\Lambda)\theta(\omega).\end{split}$ (63)
To achieve a reasonable broadening choose $A=0.025\;\mathrm{cm}^{-2}$ and $\Lambda=3500\;\mathrm{cm}^{-1}$, leading to the spectral density shown in the background of Fig. 7.
Ref. [83] shows how each of the 107 damped vibrational modes can equally be described by an effective spectral density
[…]
We thus broaden each line in Fig. 8 according to Eq. (64), which leads us to the effective spectral density seen by the electronic degree of freedom $J_{eff}(\omega)$ shown in Fig. 3(d) and the effective molecular Hamiltonian
$\begin{split}H_{mol,eff}=&\frac{\omega_{0}}{2}\sigma^{z}-\sum_{\lambda}g_{\lambda}^{eff}\sigma_{+}\sigma_{-}(b_{\lambda}+b_{\lambda}^{\dagger})\\ +&\sum_{\lambda}\omega_{\lambda}^{eff}b_{\lambda}^{\dagger}b_{\lambda},\\ J_{eff}(\omega)=&\sum_{\lambda}\lvert g_{\lambda}^{eff}\rvert^{2}\delta(\omega-\omega_{\lambda}^{eff}).\end{split}$ (65)
The spectral density $J_{eff}(\omega)$ was used to parameterize our BBGKY-HEOM model. In Fig. 8 we compare the resulting absorption spectrum of a single molecule to the path-integral implementation of ORCA [105] and observe qualitative agreement, with some quantitative deviations. A quantitative comparison is complicated by the different treatment of linewidth-broadening in ORCA.
Figure 8: Absorption spectrum of the molecule based on the Hamiltonian (65) (blue) compared to the absorption spectrum predicted by ORCA (dashed black). A quantitative comparison is complicated by the different treatment of linewidth-broadening in ORCA.
F.3 Interaction picture dynamics for few emitter lasing
From the description of a single molecule in Eq. (65) we now move to multiple molecules that couple identically with strength $g^{cav}$ to a common cavity mode that is in resonance with the electronic transition and are subject to a coherent drive with amplitude $2E_{d}$ and frequency $\omega_{0}$
$H_{tot}=$ $\sum_{j}^{N}H_{mol,eff}^{(j)}+\sum_{j}^{N}2E_{d}\cos(\omega_{0}t)\sigma_{x}^{j}$
$+\sum_{j}g^{cav}\sigma_{x}^{j}(a+a^{\dagger})+\omega_{0}a^{\dagger}a,$ (66)
where $H_{mol,eff}^{(j)}$ denotes $H_{mol,eff}$ from Eq. (65) acting on molecule $j$. Cavity loss and spontaneous decay of the individual molecules are accounted for via a master equation description, such that the total system $\rho_{tot}$ consisting of the molecules and the cavity mode evolves according to
$\dot{\rho}_{tot}=-i[H_{tot},\rho_{tot}]+\kappa\mathcal{D}[a](\rho_{tot})+\Gamma_{\downarrow}\sum_{i}\mathcal{D}[\sigma_{-}^{i}](\rho_{tot}).$ (67)
We now go into an interaction picture $\rho_{tot}^{I}=U_{I}\rho U_{I}^{\dagger}$ with
$U_{I}=\exp\left(i\omega_{0}t\left(\sum_{j,\lambda}\sigma_{z}^{j}/2+a_{\lambda}^{\dagger}a_{\lambda}\right)\right).$ (68)
After applying the rotating wave approximation as described in Sec. III.1 we end up with Eq. (III.2),(20).
Appendix G Vibrational Resonances
In the following we provide a heuristic explanation of the maxima observed in Fig. 3. A more detailed analytical investigation based on the full model instead of the toy model ind Eq. (22) is left for future work.
Starting from $H_{simple}$ in Eq. (22) we go into an interaction picture with respect to the drive |$\ket{\widetilde{\psi}}=\exp\left(i\Omega\sigma_{x}+2\Omega b^{\dagger}b\right)\ket{\psi}$ which yields
$\begin{split}\widetilde{H}_{simple}=&(\omega^{vib}-2E_{d})b^{\dagger}b+g^{vib}(\sigma_{+x}b+\sigma_{-x}b^{\dagger})\\ &+\frac{g^{vib}}{2}\left(be^{-2iE_{d}t}+b^{\dagger}e^{2iE_{d}t}\right)\\ &+\frac{g^{vib}}{2}\left(e^{-4iE_{d}t}b^{\dagger}\sigma_{+x}+e^{4iE_{d}t}b\sigma_{-x}\right),\\ \sigma_{+x}=&\frac{\sigma_{z}-i\sigma_{y}}{2},\\ \sigma_{-x}=&\frac{\sigma_{z}+i\sigma_{y}}{2}.\end{split}$ (69)
To obtain an estimate of the phonon population we neglect the time-dependent terms and obtain the dynamics of the expectation value $\langle b\rangle$ in a Heisenberg picture. We find
$\dot{\langle b\rangle}=-i(\omega^{vib}-2E_{d})\langle b\rangle-ig^{vib}\langle\sigma_{-x}\rangle.$ (70)
In the steady state we thus find $\langle b\rangle\propto\frac{1}{\omega^{vib}-2E_{d}}$. As an additional proof of this resonance phenomena is shown in Fig. 9. There we artificially modify the effective spectral density by omitting the small peak at $\omega\approx 1.3E_{max}$. We then find that also the 2nd maxima in the cavity occupation at $E_{d}\approx 0.65E_{max}$ vanishes.
Figure 9: The upper plot shows the obtained steady state cavity occupations for $N=5$ and a modified effective spectral density $J_{eff}(\omega)$, where the small peak at $\omega\approx 1.3E_{max}$ was intentionally omitted as shown in the lower plot. All other parameters are identical to Fig. 3. Compared to the results of the full model the second maxima in the cavity occupation is missing (upper plot).
Appendix H Numerical details
@@
Numerical simulations were performed with the python packages numpy[106], scipy[107] and qutip[108]. The BBGKY-HEOM equation (II.2) with the approximation for the three-body reduced density matrix given in Eq. (II.2) were solved with the scipy implementation of the explicit Runge-Kutta method of order 5(4), with an absolute tolerance of $101e-$ 10 and dynamical timestepping. We checked for convergence with respect to the hierarchy depth, which justifies a hierarchy depth of 5 for the Tavis-Cummings model and 3 for the Cavity-Fermi-Hubbard model. The benchmark simulations for the quantum master equation were performed using qutips mesolve function with default parameters.